OpenStudy (anonymous):

what is the simplified form of 3√ 5c x √ 15c^3

OpenStudy (anonymous):

3√5cx√15c3 Simplifies to: 25.980762c4x =25.980762c4x

OpenStudy (anonymous):

Its simplifying radicals

OpenStudy (anonymous):

Okay.

OpenStudy (superman499):

what do u think it is

OpenStudy (anonymous):

I'm not sure, \[3\sqrt{5c} \times \sqrt{15c _{}^{3}} \]

OpenStudy (anonymous):

then it would simply to \[3\sqrt{75\times c \times c ^{3}}\]

OpenStudy (anonymous):

actually no lol im not sure

OpenStudy (radar):

@December25 You were on the correct path:\[3\sqrt{75 \times c ^{4}}\]just continue with further simplificagtion.

OpenStudy (anonymous):

thanks i got it now

OpenStudy (radar):

Factor the 75 so that it is a product of a perfect square and an integer. Then extract square roots.

OpenStudy (radar):

Show you work and I will critique.

OpenStudy (anonymous):

\[3\sqrt{25\times 3 \times c ^{4}}\]

OpenStudy (anonymous):

\[\sqrt{25c ^{4}} \times \sqrt{3}\]

OpenStudy (anonymous):

\[15c ^{^{2}} \sqrt{3}\]

OpenStudy (radar):

Very good.